OpenStudy (anonymous):

Find the critical numbers of f, describe the open intervals on which f is increasing or decreasing, and locate all relative extrema. f(x)= x+3/x^3

OpenStudy (anonymous):

How do I solve this? I got f'(x) = 1/x^3 - 3(x+3) / x^4

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

I found -6 as a critical number .. I just need to make sure it is correct...

OpenStudy (anonymous):

intervals and extrema? is what I haven't found

jigglypuff314 (jigglypuff314):

critical number is the derivative's zero right? http://www.wolframalpha.com/input/?i=0%3D1%2Fx%5E3+-+3%28x%2B3%29+%2F+x%5E4 which would be x = -9/2

OpenStudy (anonymous):

I put in critical numbers for the derivative and it gave me -6

jigglypuff314 (jigglypuff314):

you want the critical point of f(x) you entered in the critical point for f'(x)

OpenStudy (anonymous):

it's not the critical point for the derivative?

jigglypuff314 (jigglypuff314):

the critical point of the derivative would be the original function's point of inflection. aren't you looking for the critical point for the original function?

OpenStudy (anonymous):

yeah you're right.

OpenStudy (anonymous):

is that correct for relative extrema

jigglypuff314 (jigglypuff314):

didn't the original function have a x^3 in the denominator? http://www.wolframalpha.com/input/?i=relative+extrema+of+f%28x%29+%3D+%28x%2B3%29%2Fx%5E3

OpenStudy (anonymous):

Oh my gosh. I'm sorry :( I posted it wrong no wonder we were looking at diferent things

OpenStudy (anonymous):

it's x^2.

OpenStudy (anonymous):

so the critical point would be -6

jigglypuff314 (jigglypuff314):

lol ok :) it's fine and yep :)

OpenStudy (anonymous):

and how do I get the intervals?

jigglypuff314 (jigglypuff314):

to find where the original is increasing or decreasing, find where the derivative is positive or negative

OpenStudy (anonymous):

ok my derivative is f'(x) = 1/x^2 - (2(x+3) )/ x^3

OpenStudy (anonymous):

how do I find positive and negative?

jigglypuff314 (jigglypuff314):

which can be simplified to (-x-6)/x^3 then we know that there is critical point at x=-6 and asympote at x=0 so plug in a number less then -6 into the derivative and figure out if it's positive or negative. then plug in a number between -6 and 0 into the derivative and figure out if it's positive or negative etc.

OpenStudy (anonymous):

-5 gives me 0.008

jigglypuff314 (jigglypuff314):

or you can use graph of the derivative http://www.wolframalpha.com/input/?i=0%3D1%2Fx%5E2+-+%282%28x%2B3%29+%29%2F+x%5E3 and see that before x=-6 y values are negative between -6 < x < 0 y values are positive

jigglypuff314 (jigglypuff314):

and yup :) so between -6 < x< 0 it is positive

OpenStudy (anonymous):

-7 gives me -002

OpenStudy (anonymous):

i meant .002

OpenStudy (anonymous):

and x<-6<0 is negative

OpenStudy (anonymous):

negative is maximum and positive is minimum correct?

jigglypuff314 (jigglypuff314):

negative means original function is decreasing positive means original function is increasing

jigglypuff314 (jigglypuff314):



OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

thank you

jigglypuff314 (jigglypuff314):

you're welcome :)